Determine whether the vectors are linearly dependent or independent (1,2,1),(-1,0,1) and (2,-1,4)


1
Expert's answer
2021-08-24T07:50:30-0400

"\text{Let $c_1$, $c_2$ and $c_3$ be scalars such that}
\\c_1(1,2,1) + c_2(-1,0,1) +c_3(2,-1,4) = 0
\\\text{which can be represented in matrix form as represented as}
\\\begin{vmatrix}
1 & -1 & 2\\
2 & 0 & -1\\
1 & 1 &4
\end{vmatrix}
\begin{vmatrix}
c_1\\c_2\\c_3
\end{vmatrix}
=\begin{vmatrix}
0\\0\\0
\end{vmatrix}
\\\text{Next, we calculate the determinant of the co-efficient matrix}
\\1\begin{vmatrix}
0&-1\\
1&4
\end{vmatrix}
+1\begin{vmatrix}
2&-1\\
1&4
\end{vmatrix}
+2\begin{vmatrix}
2&0\\
1&1
\end{vmatrix}
\\=1+9+4=14
\\\text{Since the determinant is not equal to zero, then the solution exists and}
\\\text{$c_1=c_2=c_3=0$}
\\\text{Therefore the vectors are linearly independent}"